OpenStudy (anonymous):

Can someone please help with this problem. A firm’s current balance sheet is as follows: Assets $100 Debt $10 Equity $90 a. What is the firm’s weighted-average cost of capital at various combinations of debt and equity, given the following information? Debit/Assets 0%, 10%, 20%, 30%, 40% 50%, 60% After-tax cost of Debt 8%, 8%, 8%, 9%, 10%, 12% Cost of Equity 12%, 12%, 12%, 13%, 14%, 15%, 16% B) Construct a pro forma balance sheet that indicates the firm's optimal capital structure. Compare this balance sheet with firm's current balance sheet. What course of action should the firm tak

OpenStudy (anonymous):

wacc=cost of debt*(portion of debt in capital structure)+cost of equity *(portion of equity in capital) if debt=10 and equity=90 cost of debt=8% cost of equity=12% (as the present ratio shows debt/asset=10/100=10% so corresponding values of cost of debt(which has no effect of tax as it is paid before tax is paid)=8%,cost of equity=12% wacc=8%*.1+12%*.9=11.6%